OpenStudy (anonymous):

Need help on rectangular distribution A random variable, X, has the pdf: f(x) = 0.05, 0

OpenStudy (anonymous):

\[0 \le x \le 20\]

OpenStudy (anonymous):

instead of 0 < x < 20

OpenStudy (zarkon):

< or \(\le\)...it doesn't matter

OpenStudy (anonymous):

I've drawn the rectangular distribution graph but got the answer to the first one wrong

OpenStudy (zarkon):

what did you get for an answer?

OpenStudy (anonymous):

Sorry, I meant I got the second one wrong

OpenStudy (anonymous):

Just realized I got the first one right

OpenStudy (zarkon):

if you can do (a) you should be able to do (b)

OpenStudy (anonymous):

I got 0.61 for some reason for b

OpenStudy (zarkon):

and how did you get that?

OpenStudy (anonymous):

20-7.8/20-0

OpenStudy (zarkon):

7.8/20

OpenStudy (anonymous):

Oh is it 7.8-0/20-0?

OpenStudy (zarkon):

you should use parentheses

OpenStudy (anonymous):

For what?

OpenStudy (anonymous):

So like (7.8-0)/(20-0)?

OpenStudy (zarkon):

7.8-0/20-0=7.8-0-0=7.8 (7.8-0)/(20-0)=\(\dfrac{7.8}{20}=.39\)

OpenStudy (anonymous):

I see, so how do I go about doing the 3c?

OpenStudy (zarkon):

\[P(a<x<b)=\frac{b-a}{20}\] for \(0\le a,b\le 20\)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I got it right

OpenStudy (zarkon):

good